OpenStudy (anonymous):

p(x)=ax²+(a-15)x+1 and p(x)=2x²-3x+1/b have real coefficients and the same roots. What's the value of a + b? a)12 b)9 c)3 d)6 How can I start solving this?

OpenStudy (bahrom7893):

Okay, if they have the same roots, the discriminant, which equals: b^2-4ac, is 0, so start off with: b^2-4ac=0

OpenStudy (bahrom7893):

oh wait nevermind that haha.. but still find the discriminant

OpenStudy (bahrom7893):

i just saw that u had two equations, i misread the question..

OpenStudy (anonymous):

oops, second one is q(x)=2x²-3x+1/b

OpenStudy (bahrom7893):

Okay, what is x equal to? x1 = (-b - sqrt(b^2-4ac))/(2a) x2 = (-b + sqrt(b^2-4ac))/(2a) right?

OpenStudy (bahrom7893):

okay let's just look at x1 for now. x1 from the first equation must equal to x1 from the 2nd one right.. so: (-b - sqrt(b^2-4ac))/(2a) = (-b - sqrt(b^2-4ac))/(2a) and if we rewrite this in terms of things given..

OpenStudy (bahrom7893):

1) First equation: a = a; b = a-15; c = 1 2) Second equation: a = 2; b = -3; c = 1/b

OpenStudy (bahrom7893):

Plug those all into the equation that i wrote above: ( -(a-15) - sqrt[(a-15)^2-4*a*1] )/(2a) = ( 3 - sqrt[9 - 4 * 2 * (1/b)] )/4

OpenStudy (bahrom7893):

Simplify the mess a little: ( -a + 15 - sqrt[ (a-15)^2 - 4a] )/(2a) = ( 3 - sqrt[9-8/b] )/4

OpenStudy (bahrom7893):

oh god.. now we gotta do the same thing for x2..

OpenStudy (phi):

Start with p(x)=ax²+(a-15)x+1 and p(x)=2x²-3x+1/b have real coefficients and the same roots. The same roots means ax²+(a-15)x+1 =0 2x²-3x+1/b = 0 for the 2nd equation, multiply both sides (all terms) by b to get 2b x^2 -3b +1 =0 and of course, ax²+(a-15)x+1 =0 now match coefficients of each term: a= 2b (a-15)= -3b solve for a and b. One way: sub in 2b for a in the 2nd equation 2b-15= -3b 5b=15 b=3 a=6 (a+b) = 9

OpenStudy (bahrom7893):

( -(a-15) + sqrt[(a-15)^2-4*a*1] )/(2a) = ( 3 + sqrt[9 - 4 * 2 * (1/b)] )/4

OpenStudy (bahrom7893):

omg phi.... lol im doing this whole freakin thing like an idiot.. fml

OpenStudy (bahrom7893):

Phi's right, all u gotta do is just match the coefficients..

OpenStudy (phi):

* 2b x^2 -3b x +1 =0 (in case it wasn't obvious)

OpenStudy (anonymous):

thank you guys very much! haha, i told ya bahrom, you like the hardest ways :P

OpenStudy (bahrom7893):

lol melinda.. that got me into trouble soo many times... but i always find the easy way at the end, like 5 mins before the exam's over